OpenStudy (anonymous):

I need to determine the indicial equation of the following equation and the exponents at the singularity for each singular point. I am starting with 0 since its a regular singular point sin(x)y''+xy'-2y=0

OpenStudy (amistre64):

"indicial equation" ?

OpenStudy (amistre64):

power series solution

OpenStudy (amistre64):

\[y=\sum_{0}a_nx^n\] \[y'=\sum_{0}a_nnx^{n-1}\] \[y''=\sum_{0}a_nn(n-1)x^{n-2}\] \[sin(x)=\sum_0\frac{(-1)^n}{(2n+1)~}x^{2n+1}\] is that sine right?

OpenStudy (amistre64):

might just wanna keep it as sin(x) for simplicity ... unless you like to play with that type of thing

OpenStudy (anonymous):

I tried putting sine as series but I don't know how to deal with it

OpenStudy (amistre64):

me either at the moment, never had the need to try to work it out :/

OpenStudy (anonymous):

I think in this case we put \[y=\Sigma_{n=0}^\infty a_nX^{r+n}\]

OpenStudy (amistre64):

lets try it as sin(x) and see where we end up subbing into the ys into the setup we get \[sin(x)\sum_0a_nn(n-1)x^{n-2}+x\sum_0a_nnx^{n-1}-2\sum_0 a_nx^n=0\] \[\sum_0sin(x)a_nn(n-1)x^{n-2}+\sum_0a_nnx^{n}+\sum_0 -2a_nx^n=0\] \[\sum_0sin(x)a_nn(n-1)x^{n-2}+\sum_0a_n(n-2)x^{n}=0\] let get the exponents to even up, lowest rules \[\sum_0sin(x)a_nn(n-1)x^{n-2}+\sum_{0+2}a_{n-2}(n-2-2)x^{n-2}=0\] \[\sum_0sin(x)a_nn(n-1)x^{n-2}+\sum_{2}a_{n-2}(n-2-2)x^{n-2}=0\] and just line up our indexes

OpenStudy (amistre64):

\[sin(0)a_0~0(0-1)x^{0-2}\\+sin(1)a_1~1(1-1)x^{0-2}\\+\sum_2sin(x)a_nn(n-1)x^{n-2}+\sum_{2}a_{n-2}(n-2-2)x^{n-2}=0\] \[sin(0)a_0~0(0-1)x^{0-2}\\+sin(1)a_1~1(1-1)x^{1-2}\\+\sum_2sin(x)a_nn(n-1)x^{n-2}+\sum_{2}a_{n-2}(n-4)x^{n-2}=0\] or does heighest exponent rule?

OpenStudy (amistre64):

either way it seems to pan out the same on that

OpenStudy (amistre64):

the key is when all a_n parts are zero

OpenStudy (amistre64):

\[\sum_2sin(x)a_nn(n-1)x^{n-2}+\sum_{2}a_{n-2}(n-4)x^{n-2}=0~:~a_0=0,a_1=0\] \[\sum_2[sin(x)a_nn(n-1)+a_{n-2}(n-4)]x^{n-2}=0\] \[sin(x)a_nn(n-1)+a_{n-2}(n-4)=0\] \[sin(x)a_nn(n-1)=-a_{n-2}(n-4)\] \[a_n=-a_{n-2}\frac{(n-4)}{n(n-1)sin(x)}\]

OpenStudy (amistre64):

i might wanna solve that for an-2 instead of an

OpenStudy (anonymous):

if we put it in the equation we would get:, I did \[\sum_0 \frac{(-1)^n X^{2n+1}}{(2n+1)!} \sum_{0}(r+n)(r+n-1)a_nX^{r+n-2} +\sum_{0}(r+n)X^{r+n}-\sum_0 2a_n X^{r+n}\] then change the exponent t \[\sum_0 \frac{(-1)^n X^{2n+1}}{(2n+1)!} \sum_{-2}(r+n+2)(r+n+1)a_{n+2}X^{r+n} +\sum_{0}(r+n)X^{r+n}-\sum_0 2a_n X^{r+n}\] but then I don't know what to do I was thinking of expandind the first few terms at a0 \[[1-\frac{X^3}{3!}..]*[(r(r-1)a_0X^{r-2}+...]+[ra_oX^r+...]-2a_0X^r+...] \] and combine??

OpenStudy (anonymous):

because it is't the recurrence relation that we need , we want the indicial equation

OpenStudy (amistre64):

\[a_2=a_{0}\frac{4-2}{2(1)sin(x)}\] \[a_4=a_{2}\frac{4-4}{4(3)sin(x)}\\~~~~~~=a_{0}\frac{4-2}{2(1)sin(x)}\frac{4-4}{4(3)sin(x)}\] and the recurrence equation unfolds as such

OpenStudy (amistre64):

i cant say im all that familiar with your notational setup

OpenStudy (anonymous):

we are suppose to get a characteristic equation such as r(r-1) =0 and then get a 2 exponents which we replace r and then find the soulution as we would do normally

OpenStudy (anonymous):

this explains it better http://www.math.mcgill.ca/jakobson/courses/ma261/lecture17.pdf

OpenStudy (amistre64):

i can make sense of about half of that :)

OpenStudy (amistre64):

the product of sums still gets me

OpenStudy (amistre64):

using some simple summations, how would you go about explaining the process? or do you only know how the Frobe is spose to look?